OpenStudy (anonymous):

3/7 x 5/6 = 3/7 x 5/6 = 15/42 divided by 3/3 = 5/14 is this true?

OpenStudy (anonymous):

Why are u posting the question in 2places at once?

jimthompson5910 (jim_thompson5910):

This is correct \[\frac{3}{7}\times\frac{5}{6}=\frac{5}{14}\]

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

but how do i get 15/42 divided by 3/3

OpenStudy (anonymous):

because 3 is a common factor